OpenStudy (anonymous):

How do you find the common denominator for 80 and 1260/23

OpenStudy (anonymous):

do you mean lcd for 80 and 23?

OpenStudy (anonymous):

I am not sure, I will just give you the whole problem k solve using elimination

OpenStudy (anonymous):

0.3x-0.2y=4 0.2x+0.5y=-21/23

OpenStudy (anonymous):

just a quick answer as i must go multiply first equation by 2 and second by 3 then subtract them to eliminate x solve resulting equation for y then plug value of y into one of the equations to find value of x must go

OpenStudy (anonymous):

I multiplyed by 10 to get rid of decimals 3x-2y=40 2x+5y=-210/23 then multiply the first equation by 2 and the second by -3 6x-4y=80 -6x+-15y=

OpenStudy (anonymous):

okay thanks I will try that

OpenStudy (amistre64):

80/1 and 210/23 ; have 23 as a common denominator. multiply 80/1 by a useful for of 1 in order to change the way it looks without changeing its value. \[\frac{80}{1}*\frac{23}{23}=\frac{80(23)}{23}\]